please help. thank you so much !!
0.0702 0.0702 incorrect incorrect At least one of the answers above is NOT correct. 1 of the questions remains unanswered. (1 1
0.0201 0.0201 incorrect A correct At least one of the answers above is NOT correct. (1 point) 35 people are randomly selected2
Answer #1

Son: Giuen, =100 68.31 = 91.5 - To test Ho el=go Us HI; +90 Test statistic; 3-il 11 91.5-90 8.31 / 100 1.5 N 11 0.831 Z - 1.8La 2) 35, 0.02 = 99 sec : S228 sec To test. Ho ll= o see NS Hi;el to sec Test statistic 7 = X-el San 94-0 8 229 / 135 2.44 で